OpenStudy (anonymous):

write y=1/8x+7 in standard form using integers. something about multiplying 7 to both sides..?

OpenStudy (anonymous):

If you want it in standard form, this is what it should look like: \[\large ax+by+c=0\]

OpenStudy (anonymous):

get rid of the fraction you have by multiplying both sides by the number on that denominator of the fraction.

OpenStudy (anonymous):

how do i make it into that form then? there's an equation to do that right?

OpenStudy (anonymous):

what do you mean both sides? 7 and what? 1 or 1/8?

OpenStudy (anonymous):

Look at the fraction that you have in your equation.

OpenStudy (anonymous):

The standard form is ax+by=c. so, y=1/8x+7. You're trying to make the equation to ax+by. so, the x and the y need to be in the same side. To do that, you need to add y to the x side. So the equation is going to be 1/8+7-y=0. Then, you subtract 7 from both side, which gives you 1/8x-y=-7.

OpenStudy (anonymous):

All you have to do now @sohostarlight is to get rid of the fraction. To do that, you multiply both sides by the number on the denominator.

OpenStudy (anonymous):

-x+8y=56?

OpenStudy (anonymous):

@Azteck @Xeos

OpenStudy (anonymous):

Yes, it could be like that.

OpenStudy (anonymous):

Incorrect.

OpenStudy (anonymous):

It should be recommended that the x be positive.

OpenStudy (anonymous):

If you give the signs a switcheroo, then you're done.

OpenStudy (anonymous):

@Azteck so its -x+8y=56?